# ModServe: Scalable and Resource-Efficient Large Multimodal Model Serving

### v2

@@ Abstract.
Large multimodal models (LMMs) demonstrate impressive capabilities in understanding images, videos, and audio beyond text. However, efficiently serving LMMs in production environments poses significant challenges due to their complex architectures and heterogeneous characteristics across their multi-stage inference pipelines.
We present the first comprehensive systems analysis of two prominent LMM architectures, decoder-only and cross-attention, across six representative open-source models, revealing key systems design implications. We also present an in-depth analysis of production LMM inference traces, uncovering unique workload characteristics, including variable, heavy-tailed request distributions and bursty traffic patterns.
Based on these insights, we propose ModServe, a modular LMM serving system that decouples stages for independent optimization and adaptive scaling. ModServe dynamically reconfigures stages and handles bursty traffic with modality-aware scheduling and autoscaling to meet tail latency SLOs while minimizing costs. ModServe achieves 3.3–5.5 higher throughput (leading to 25–41.3% cost saving) while meeting SLOs on a 128-GPU cluster with production traces.
@@ 1. Introduction
The rapid advancement in generative AI has led to the development of large multimodal models (LMMs) capable of processing inputs across various modalities such as text, image, video, and audio. These models have demonstrated remarkable capabilities in tasks like image captioning (Chen et al., 2022; Mokady et al., 2021; Hu et al., 2023), visual question answering (Schwenk et al., 2022; Shao et al., 2023), and multimodal dialogue systems (Li et al., 2024b; Chen et al., 2024b; Team et al., 2024). This has led to a rapid adoption of LMMs in production services, including online applications where latency service-level objectives (SLOs) are critical.
@@
Unlike traditional large language models (LLMs) that process purely textual inputs using a single component, a decoder-based transformer architecture (Waswani et al., 2017), LMMs handle fundamentally different types of inputs, each requiring distinct processing approaches. This heterogeneity introduces unique serving complexities that demand novel analysis and serving strategies. For Image-Text-to-Text models (HuggingFace, 2024c), the inference pipeline consists of multiple specialized stages: image preprocessing to transform raw images into tensor representations, image encoding to convert these tensors into image tokens, and a language model backend that combines text prompts with image tokens to generate text outputs. Currently, these stages are typically served as a monolithic system (Kwon et al., 2023; Wolf et al., 2020; Aminabadi et al., 2022), where all components are integrated within a single serving instance and scaled together as a unified entity.
@@
Unfortunately, existing monolithic inference serving systems fail to account for multimodality, making them unable to scale effectively while meeting time-to-first-token (TTFT) SLOs, which now include image processing and encoding times. Figure 1 shows how a monolithic deployment struggles to scale as the number of images per request increases (a common scenario in multi-image or video workloads) resulting in sharp TTFT degradation. As a result, image-heavy requests can result in head-of-line (HoL) blocking, reducing system responsiveness and causing overprovisioning.
@@
Our Work. In this paper, we present the first comprehensive systems analysis of two leading LMM architectures: cross-attention (CroAttn) and decoder-only (DecOnly), on both open-source LMMs and novel production LMM inference traces in Azure datacenters. We analyze their multi-stage inference pipelines, performance-resource tradeoffs, and production workload patterns, including variable request rates, diverse multimodal inputs, and bursty traffic. We focus on Image-Text-to-Text but our insights extend to other multimodal scenarios, such as Video-Text-to-Text, where videos are processed as image frame sequences (Li et al., 2024c), and Audio-Text-to-Text tasks (HuggingFace, 2024a), which share similar model architectures with the models we study.
@@
Based on these insights, we propose ModServe, a novel modular architecture for scalable and resource-efficient LMM serving which directly addresses the challenges identified in our analysis. ModServe separates image- and text-specific inference stages into distinct instances for decoupled execution. In ModServe, Image Instances handle image preprocessing and encoding, while Text Instances manage LLM prefill and decoding (Figure 1). Text-only requests are served by Text Instances, whereas image-text requests go through Image Instances where images are converted to tokens before being forwarded to Text Instances for text generation.
@@
Further, to manage image-driven bursts, ModServe implements modality-aware routing for Image and Text Instances. For example, images from image-text requests are routed to Image Instances with the fewest pending image tokens to encode, reducing HoL blocking and tail latency spikes.
We implement ModServe on top of a high-performance inference system, vLLM (Kwon et al., 2023), and demonstrate the effectiveness of ModServe through extensive evaluations on a 16-server (128 GPUs) cluster running production Azure LMM inference traces. Compared to state-of-the-art baselines, ModServe achieves 3.3–5.5 higher throughput under static allocation and reduces LMM serving cost by 25–41.3% while meeting the P99 TTFT SLOs.
@@
A comprehensive system characterization on LMM serving, examining performance profiles and resource utilization patterns across diverse workloads in both open-source LMM deployments and production environment.
A large open-source dataset containing sampled production Azure LMM inference traces.
Design and implementation of ModServe, a modular architecture for scalable and efficient LMM serving.
@@ 2. Large Multimodal Models Background
| (#Params) | () | (#Params) | Parallelism | (HF-VLM (Duan et al., 2024)) | ||||
@@
Decoder-Only (DecOnly) LMMs. An unmodified LLM backend is reused in DecOnly LMMs (e.g., LLaVA-OV reuses Qwen2 LLM (Li et al., 2024c)), processing text and image tokens uniformly (“DecOnly” box in Figure 2). While valued for their simplicity and unified modality handling, DecOnly models often require long sequences for high-resolution images, resulting in computational inefficiencies during inference.
@@
SLO Metrics for LMM Inference. Production LMM serving systems need to satisfy SLOs defined on tail latency (e.g., P99) for worst-case performance. These SLO metrics include Time to First Token (TTFT) and Time Between Tokens (TBT). TTFT measures the latency from query (text/images) to the first response token, critical for interactive applications. In contrast to text-centric LLM serving, LMM TTFT includes the following stages of LMMs inference pipeline: (1) image preprocessing, (2) encoding, and (3) language model prefill time. TBT captures the delay between consecutive token generations during decoding, affecting output fluency. As multimodal preprocessing and encoding primarily influence TTFT, in this work, we focus on TTFT while leveraging state-of-the-art techniques (Agrawal et al., 2024; Patel et al., 2024) to optimize TBT.
An ideal LMM-serving system should meet TTFT/TBT SLOs while maximizing request throughput (i.e., goodput) and compute utilization (GPU cost).
LMM Deployments Today. State-of-the-art serving frameworks (Kwon et al., 2023; Wolf et al., 2020; Aminabadi et al., 2022) deploy LMMs as monolithic systems to meet latency SLOs. In this setup, all inference components (i.e., image preprocessor, image encoder, and LLM backend) are co-located on the same hardware server as a single unit. These tightly coupled components share uniform batching and model parallelism strategies across the pipeline. Table 1 details the default model parallelism for our open-source LMMs. While this monolithic design is straightforward to implement and common in open-source LMM serving, it limits flexibility and suffers from sharp TTFT degradation under image-heavy workloads (Figure 1).
@@ 3. Motivation and LMM Characterization
To further understand the limitations of monolithic deployments and explore unique characteristics that distinguish LMM serving from text-centric LLM serving, we characterize open-source LMMs in the Image-Text-to-Text category (HuggingFace, 2024c). We evaluate the performance and resource characteristics of heterogeneous inference stages under varying image inputs and model configurations (Section 3.1). Moreover, to understand multimodal traffic patterns at scale, we analyze sample production traces from one production LMM inference cluster in Azure (Section 3.2).
@@
Models. We use six representative open-source LMMs across two different architectures (DecOnly and CroAttn) as listed in Table 1. We deploy the models on vLLM (Kwon et al., 2023) in BF16.
@@ 3.1. Characterization on Open-Source LMMs
Per-stage Latency Breakdown. Figure 5 plots the split-up of TTFT across the three stages that comprise it; image preprocessing, image encoding, and LLM prefill. There are three key takeaways. First, image preprocessing, which occurs on the CPU, contributes minimally to the overall TTFT, while image encoding time contributes to a major portion in TTFT (especially for CroAttn models). For instance, 79% and 65% of TTFT in Llama3.2-11B and Llama3.2-90B are from image encoding. For DecOnly models such as InternVL-26B and NVLM-D-72B, image encoding latency accounts for 25% and 54% of TTFT. Second, the image encoding time depends on the encoder model size. For instance, scaling from SigLIP-400M (in LLaVA-OV-7B) to InternViT-6B (in InternVL-26B), the median image encoding time increases by 10. Finally, prefill computation is more efficient in CroAttn models because image tokens are attended to only in the CroAttn layers, as previously discussed in Section 2.
Compute Characteristics of LMM Stages. Image preprocessing on CPU and image encoding on GPU are compute-intensive processes. Figure 6(a) plots the impact of varying the number of CPU cores on preprocessing latency. Preprocessing is CPU-intensive and benefits from trivially parallelizing across cores. Both stages exhibit linear latency scaling with batch size, saturating compute without significant throughput gains from increased batching as shown in Figures 6(b) and 6(c), respectively. Figure 6(d) further plots the GPU utilization metrics for a request batch size of one during image preprocessing and image encoding. We observe a consistent SM core activity near 100% during image encoding, with average DRAM utilization below 30%. Image encoding is, therefore, typically compute-bound, resembling the language model’s prefill phase (Kamath et al., 2024). Moreover, when a request has multiple images in the input prompt, there is no compute dependency between the images during image encoding; hence, image tiles can be parallelized across multiple encoders.
Compute characteristics of prefill and decode phases of the LLM backend have been well studied; the prefill phase is typically compute-bound, while the decode phase is memory-bound (Patel et al., 2024; Agrawal et al., 2024; Kamath et al., 2024). However, Figure 5 shows that LLM prefill is more efficient in CroAttn models than inDecOnly models, resulting in reduced compute boundness and an interesting tradeoff we describe below.
@@
Impact of Parallelism. Monolith deployments also limit the flexibility of model sharding within a GPU server which is typically done through tensor parallelism (TP) (vLLM, 2024). Figure 9 shows how increasing TP degrees affects latency across LMM stages. In Llama3.2-11B, the lowest LLM prefill latency occurs at TP-8, image encoding at TP-4, and TBT at TP-1. At TP-8, encoding latency rises due to the tradeoff between compute intensity and inter-GPU communication, making it inefficient to split a small 630M encoder across eight GPUs.
@@ 3.2. Production LMM Trace Analysis
Building on the insights from open-source LMM characterization, we further analyze multimodal traffic patterns at scale, leveraging production traces from one of Azure’s LMM inference clusters. The traces capture a sample of multi-tenant traffic, including both text-only and image-text requests. Our study focuses on (1) temporal and burstiness patterns and (2) heterogeneity of multimodal requests. We plan to open-source these production traces.
@@
Image-Driven Bursts. Image-text requests experience significant burstiness, not only due to higher request arrival rates but also increased images per request (e.g., video workloads). As a result, existing LLM traffic prediction methods (Stojkovic et al., 2025b) (which work well for workloads with diurnal patterns) have a high average error rate of 79%.
Request Heterogeneity. Figure 11(a) shows that prompt lengths vary significantly across modalities. Both image-text and text-only requests follow a heavy-tailed power-law distribution (=4.4 and 2.9, respectively) where a higher means a heavier tail with more extreme events occurring more frequently. In addition, image-text requests have longer median prompts due to image tokens but shorter tails than text-only requests. Figure 11(b) shows that the number of images per request also varies significantly with a heavy tail. In addition, among the top three services issuing text-image requests, we observe high inter-service variability. Some services process 16 more images per request than others.
@@
TTFT increases with the percentage of image tokens in a request due to the increased image encoding computation, resulting in a 1.5 TTFT degradation when transitioning from text-only to image-only inputs. However, this latency gain is significantly lower than DecOnly models because CroAttn models attend to image tokens only within the CroAttn layers, resulting in reduced LLM prefill time (shown in green) and partially offsetting the overhead from image encoding. The right figure further illustrates this by breaking down the layer-wise LLM prefill time, highlighting a reduction in self-attention compute (i.e., “Attn Layer” and “MLP Layer”) as the proportion of image tokens increases. Although the cross-attention computation peaks at the 50% image tokens (due to the dependency on both image and text tokens), it contributes much less than self-attention computation because there are only 4 CroAttn layers (out of 40 layers).

### v3

# ModServe: Modality- and Stage-Aware Resource Disaggregation for Scalable Multimodal Model Serving

@@ Abstract.
Large multimodal models (LMMs) demonstrate impressive capabilities in understanding images, videos, and audio beyond text. However, efficiently serving LMMs in production environments poses significant challenges due to their complex model architectures and heterogeneous characteristics across their multi-stage inference pipelines and modalities.
We present the first comprehensive systems analysis of two prominent LMM architectures, decoder-only and cross-attention, across six representative open-source models, revealing key systems design implications. We also present an in-depth analysis of production LMM inference traces, uncovering unique multimodal workload characteristics, including variable, heavy-tailed request distributions and bursty traffic patterns.
Based on these insights, we propose ModServe, a modular LMM serving system that decouples model stages for independent optimization and adaptive scaling. ModServe dynamically reconfigures stages and handles bursty traffic with modality-aware scheduling and autoscaling to meet tail latency SLOs while minimizing costs. ModServe achieves 3.3–5.5 higher throughput (leading to 25–41.3% cost saving) while meeting SLOs on a 128-GPU cluster with production multimodal traces.
@@ 1. Introduction
The rapid advancement in generative AI has led to the development of large multimodal models (LMMs) capable of processing inputs across various modalities such as text, image, video, and audio. These models have demonstrated remarkable capabilities in tasks like image captioning (Chen et al., 2022; Mokady et al., 2021; Hu et al., 2023), visual question answering (Schwenk et al., 2022; Shao et al., 2023), multimodal dialogue systems (Li et al., 2024b; Chen et al., 2024b; Team et al., 2024), and voice assistant (He et al., 2025). This has led to a rapid adoption of LMMs in production services, including online user-facing applications where latency service-level objectives (SLOs) are critical.
@@
Unlike traditional large language models (LLMs) that process purely textual inputs using a single component, a decoder-based transformer architecture (Waswani et al., 2017), LMMs handle fundamentally different types of inputs, each requiring distinct processing approaches. This heterogeneity introduces unique serving complexities that demand novel analysis and serving strategies. For Image-Text-to-Text models (HuggingFace, 2024c), the inference pipeline consists of multiple specialized stages: image preprocessing to transform raw images into tensor representations, image encoding to convert these tensors into image tokens, and a language model backend that combines text prompts with image tokens to generate text outputs. Currently, these stages are typically served as a monolithic system (Kwon et al., 2023; Wolf et al., 2020; Aminabadi et al., 2022), where all components are integrated within a single serving instance and scaled together as a unified entity. While recent LLM serving systems adopt prefill-decode (PD) disaggregation (Patel et al., 2024; Zhong et al., 2024) to reduce the performance interference at the LLM backend, these optimizations remain text-centric and overlook the upstream stages of multimodal preprocessing and encoding, which are still tightly coupled within a monolithic serving instance.
@@
This lack of modality-aware decoupling limits the efficiency and scalability of existing monolithic inference serving systems when serving workloads beyond text. These systems struggle to meet time-to-first-token (TTFT) SLOs because multimodal input preprocessing and encoding are in the critical path. Figure 1 illustrates the challenges faced by a monolithic deployment in scaling as the number of images per request increases (a common scenario in multi-image or video workloads), resulting in sharp TTFT degradation. As a result, image-heavy requests can result in head-of-line (HoL) blocking, reducing system responsiveness and causing resource overprovisioning.
@@
Our Work. In this paper, we present the first comprehensive systems analysis of two leading LMM architectures: cross-attention (CroAttn) and decoder-only (DecOnly), on both open-source LMMs and novel production LMM inference traces in Azure datacenters. We analyze their multi-stage inference pipelines, performance-resource tradeoffs, and production workload patterns, including variable request rates, diverse multimodal inputs, and bursty traffic. To advance research in this area, we release the first open-source multimodal trace from our production clusters, enabling the community to study real-world deployment patterns.
@@
Based on these insights, we propose ModServe, a novel modular architecture for scalable and resource-efficient LMM serving which directly addresses the challenges identified in our analysis. Take Image-Text-to-Text tasks as an example, ModServe separates image- and text-specific inference stages into distinct instances for decoupled execution. In ModServe, Image Instances handle image preprocessing and encoding, while Text Instances manage LLM prefill and decoding (Figure 1). Text-only requests are served by Text Instances, whereas image-text requests go through Image Instances where images are converted to tokens before being forwarded to Text Instances for text generation.
@@
Further, to manage image-driven bursts, ModServe implements modality-aware request routing for Image and Text Instances. For example, images from image-text requests are routed to Image Instances with the fewest pending image tokens to encode, reducing HoL blocking and tail latency spikes.
We implement ModServe on top of a high-performance inference system, vLLM (Kwon et al., 2023), and demonstrate the effectiveness of ModServe through extensive evaluations on a 16-server (128 GPUs) cluster running production LMM inference traces from Azure. Compared to state-of-the-art baselines, ModServe achieves 3.3–5.5 higher throughput under static allocation and reduces LMM serving cost by 25–41.3% while meeting the P99 TTFT SLOs.
While existing techniques like PD disaggregation have shown promise for optimizing LLM inference by separating prefill and decode phases (Zhong et al., 2024; Patel et al., 2024), they fall short for multimodal workloads. The unique nature of bursty and heterogeneous multimodal traffic observed in production makes it nontrivial to extend PD disaggregation studied in the context of text-centric LLM inference; modality-specific optimizations such as parallelizing encoding, modality-aware routing, and stage autoscaling are crucial to meet SLOs while maintaining resource efficiency. Additionally, the choice of selecting PD disaggregation or colocation for the LLM backend depends on the workload conditions and optimization targets (Zhong et al., 2024). We show in Section 5 that ModServe is composable with both colocated LLM backends (mixed PD batching) as well as PD disaggregation for the text nodes, and improves serving latency with the proposed modality-aware techniques in both LLM backend configurations.
We focus on Image-Text-to-Text and Video-Text-to-Text (where videos are processed as image frame sequences (Li et al., 2024c)), but our insights extend to other multimodal scenarios, such as Audio-Text-to-Text tasks (HuggingFace, 2024a), which share similar model architectures and inference stages with the models we study.
@@
The first open-source dataset containing large-scale production LMM inference traces from Azure (GitHub, 2025).
A comprehensive system characterization on LMM serving, examining performance profiles and resource utilization patterns across diverse multimodal workloads in both open-source LMM deployments and production environments.
Design and implementation of ModServe, a modular architecture for scalable and resource-efficient LMM serving.
@@ 2. Large Multimodal Models Background
| (#Params) | () | (#Params) | Parallelism | (Open VLM Benchmark (Duan et al., 2024)) | ||||
@@
Decoder-Only (DecOnly) LMMs. An unmodified LLM backend is reused in DecOnly LMMs (e.g., LLaVA-OV reuses Qwen2 LLM (Li et al., 2024c)), processing text and image tokens uniformly (shown as the “DecOnly” box in Figure 2). This works by attaching a connector (Zhu et al., 2025) or modality-alignment module (e.g., MLP) that maps the image encoder output into the LLM’s token space. While valued for their simplicity and unified modality handling, DecOnly models often require long sequences for high-resolution images, resulting in computational inefficiencies during inference.
@@
SLO Metrics for LMM Inference. Production LMM serving systems need to satisfy SLOs defined on tail latency (e.g., P99) for worst-case performance. These SLO metrics include Time to First Token (TTFT) and Time Between Tokens (TBT). TTFT measures the latency from query (text/images) to the first response token, critical for interactive applications. In contrast to text-centric LLM serving, LMM TTFT includes the following stages of LMMs inference pipeline: (1) image preprocessing, (2) encoding, and (3) language model prefill time. TBT captures the delay between consecutive token generations during decoding, affecting output fluency. As multimodal preprocessing and encoding primarily influence TTFT, in this work, we focus on TTFT while leveraging state-of-the-art techniques (Agrawal et al., 2024a; Patel et al., 2024) to optimize TBT. An ideal LMM-serving system should meet TTFT/TBT SLOs while maximizing request throughput (i.e., goodput) and compute utilization (GPU cost).
LMM Deployments Today. State-of-the-art serving frameworks (Kwon et al., 2023; Wolf et al., 2020; Aminabadi et al., 2022) deploy LMMs as monolithic systems to meet latency SLOs. In this setup, all inference components (i.e., image preprocessor, image encoder, and LLM backend) are co-located on the same hardware server as a single unit. While PD disaggregation can be applied within this monolithic setup to decouple prefill from decode phases and reduce interference during text generation, it still couples multimodal components with the prefill instances (Kwon et al., 2023), preventing independent optimization of each component based on their distinct resource requirements and performance characteristics. These tightly coupled components share uniform batching and model parallelism strategies across the pipeline. Table 1 details the default model parallelism for our open-source LMMs. While this monolithic design is straightforward to implement and common in open-source LMM serving, it limits flexibility and suffers from sharp TTFT degradation under image-heavy workloads (Figure 1).
@@ 3. Motivation and LMM Characterization
To further understand the limitations of monolithic deployments and explore unique characteristics that distinguish LMM serving from text-centric LLM serving, we characterize open-source LMMs in the Image-Text-to-Text category (HuggingFace, 2024c). We evaluate the performance and resource characteristics of heterogeneous inference stages under varying image inputs and model configurations (Section 3.1). Moreover, to understand multimodal traffic patterns at scale, we analyze sample production traces from one production LMM inference cluster at Azure (Section 3.2).
@@
Models. We use six representative open-source LMMs across two different architectures (DecOnly and CroAttn) as listed in Table 1. We deploy the models on vLLM (Kwon et al., 2023) in BF16 with the default, stable PD colocated mode because the experimental PD disaggregated mode of vLLM has limited support for multimodality, and no support for cross-attention models. However, we later show in Section 5.4 that PD disaggregation is an optimization that is trivially composable with our design in ModServe.
@@ 3.1. Characterization on Open-Source LMMs
Per-stage Latency Breakdown. Figure 5 plots the split-up of TTFT across the three stages that comprise it; image preprocessing, image encoding, and LLM prefill. There are three key takeaways. First, image preprocessing, which occurs on the CPU, contributes minimally to the overall TTFT, while image encoding time contributes to a major portion in TTFT (especially for CroAttn models). For instance, 79% and 65% of TTFT in Llama3.2-11B and Llama3.2-90B are from image encoding. For DecOnly models such as InternVL-26B and NVLM-D-72B, image encoding latency accounts for 25% and 54% of TTFT. Second, the image encoding time depends on the encoder model size. For instance, scaling from SigLIP-400M (in LLaVA-OV-7B) to InternViT-6B (in InternVL-26B), the median image encoding time increases by 10. Since connectors (Zhu et al., 2025) are extremely small (e.g., of total parameters in InternVL-26B), they contribute negligible latency ( of TTFT). Finally, prefill computation is more efficient in CroAttn models because image tokens are attended to only in the CroAttn layers while the majority of LLM backends are self-attention layers, as described in Section 2.
Taken together, these results underline why image encoding emerges as a major bottleneck in multimodal model serving. First, the number of images per request can be substantial, especially in video understanding or multi-image tasks. Second, CroAttn models shift the computational load to the image encoder by reducing image-text interaction in the LLM. This lowers the computational load from LLM and keeps it in encoding.
Compute Characteristics of LMM Stages. Image preprocessing on CPU and image encoding on GPU are compute-intensive processes. Figure 6(a) plots the impact of varying the number of CPU cores on preprocessing latency. Preprocessing is CPU-intensive and benefits from trivially parallelizing across cores. Both stages exhibit linear latency scaling with batch size, saturating compute without significant throughput gains from increased batching as shown in Figures 6(b) and 6(c), respectively. Figure 6(d) further plots the GPU utilization metrics for a request batch size of one during image preprocessing and image encoding. We observe a consistent SM core activity near 100% during image encoding, with average DRAM utilization below 30%. Image encoding is, therefore, typically compute-bound, resembling the language model’s prefill phase (Kamath et al., 2024). Moreover, when a request has multiple images in the input prompt (e.g., video workloads), there is typically no compute dependency between the images during image encoding; hence, image tiles can be parallelized across multiple encoders.
Compute characteristics of prefill and decode phases of the LLM backend have been well studied; the prefill phase is typically compute-bound, while the decode phase is memory-bound (Patel et al., 2024; Agrawal et al., 2024a; Kamath et al., 2024). However, Figure 5 shows that LLM prefill is more efficient in CroAttn models than in DecOnly models, resulting in reduced compute intensity and an interesting tradeoff we describe below.
@@
Impact of Parallelism. Monolith deployments also limit the flexibility of model sharding within a GPU server which is typically done through tensor parallelism (TP) (vLLM, 2024). Figure 9 shows how increasing TP degrees affects latency across LMM stages. In Llama3.2-11B, the lowest LLM prefill latency occurs at TP-8, image encoding at TP-4, and TBT at TP-1. At TP-8, encoding latency rises due to the tradeoff between compute intensity and inter-GPU communication, making it inefficient to split a 630M-sized encoder across 8 devices.
@@ 3.2. Production LMM Trace Analysis
Building on the insights from open-source LMM characterization, we further analyze multimodal traffic patterns at scale, leveraging production traces from one of Azure’s LMM inference clusters. The traces capture a sample of multi-tenant traffic, including both text-only and image-text requests. Our study focuses on (1) temporal and burstiness patterns and (2) heterogeneity of multimodal requests. We have released these multimodal inference traces at https://github.com/Azure/AzurePublicDataset.
@@
Image-Driven Bursts. Image-text requests experience significant burstiness, not only due to higher request arrival rates but also increased images per request (e.g., video workloads). As a result, existing LLM traffic prediction methods (Stojkovic et al., 2025b) (which work well for LLM workloads with diurnal patterns) have a high average prediction error rate of 79%.
Request Heterogeneity. Figure 11(a) shows that prompt lengths vary significantly across modalities. Both image-text and text-only requests follow a heavy-tailed power-law distribution (=4.4 and 2.9, respectively) where a higher means a heavier tail with more extreme events occurring more frequently. In addition, image-text requests have longer median prompts due to image tokens but shorter tails than text-only requests. Figure 11(b) shows that the number of images per request also varies significantly with a heavy tail. In addition, among the top three services issuing text-image requests, we observe high inter-service variability. Some services (e.g., video) process 16 more images per request than others.
@@
TTFT increases with the percentage of image tokens in a request due to the increased image encoding computation, resulting in a 1.5 TTFT degradation when transitioning from text-only to image-only inputs. However, this latency gain is significantly lower than DecOnly models because CroAttn models attend to image tokens only within the CroAttn layers, resulting in reduced LLM prefill time (shown in green) and partially offsetting the overhead from image encoding. The right figure in Figure 12 further illustrates this by breaking down the layer-wise LLM prefill time, highlighting a reduction in self-attention compute (i.e., “Attn Layer” and “MLP Layer”) as the proportion of image tokens increases. Although the cross-attention computation peaks at the 50% image tokens (due to the dependency on both image and text tokens), it contributes much less than self-attention computation because there are only 4 CroAttn layers (out of 40 layers).